27 drops each of radius 0.02m carries a charge of 5 mf and combined to form a big drop compare the surface surface densities of the drop before and after merge
 

Grade:10

1 Answers

Vikas TU
7 years ago
Conserve the volume simply.
27*(4/3)*3.14*0.02^3 = (4/3)*3.14*R^3
R^3 = 3*0.02 = 0.06 m.
Now u can easily calculate their surface densities.